OpenStudy (anonymous):

Could someone help me with Radical Equations and Inequalities? One of the questions is: The solution to the radical function: 2sqrtx+5=8 - - x=11 True or False.. I said false, but I don't know if that's entirely correct...

OpenStudy (anonymous):

the equation is 2sqrt(x+5)=8 => sqrt(x+5)=8/2=4 => x+5=4^2=16 => x=16-5=11